# Optimal Equivariant Matchings on the 6-Cube With an Application to the King Wen Sequence

### v1

@@ 1 Introduction
The Boolean hypercube $\{0,1\}^{n}=\{0,1\}^{n}$ admits a natural action by the Klein four-group $K_{4}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, generated by bitwise complement $\mathrm{comp}$ and bit reversal $\mathrm{rev}$. A $K_{4}$-equivariant perfect matching is a partition of $\{0,1\}^{n}$ into pairs such that paired elements lie in the same $K_{4}$-orbit.
This paper addresses the following optimization problem:
Among $K_{4}$-equivariant perfect matchings on $\{0,1\}^{n}$ where each pair uses either $\mathrm{comp}$ or $\mathrm{rev}$, which minimize total Hamming distance (Definition 2.3)?
For $n=6$, we prove uniqueness: there is exactly one optimal matching of this form, determined by a simple greedy rule. As an application, we verify that the King Wen sequence—a traditional ordering of the 64 hexagrams of the I Ching into 32 pairs—is isomorphic to this unique optimal matching. We also note that relaxing the constraint to allow $\mathrm{comp}\circ\mathrm{rev}$ pairings yields a lower-cost matching (96 vs 120), but requires per-orbit case analysis rather than the uniform reverse-priority rule.
@@ 1.1 Main Results
The King Wen sequence of the I Ching defines a perfect matching on $\{0,1\}^{6}$ that is isomorphic to the unique optimal $K_{4}$-equivariant matching of Theorem 1.1.
@@ 2.1 Hexagrams as Binary Vectors
A hexagram is an element $h\in\{0,1\}^{6}=\{0,1\}^{6}$. We index positions 0 (bottom) through $5$ (top), following the traditional convention that hexagrams are read from bottom to top. The correspondence between hexagrams and binary numbers was noted by Leibniz [2].
There are $2^{6}=64$ hexagrams, corresponding to the vertices of the 6-dimensional hypercube.
@@ 2.2 The Klein Four-Group Action
For any hexagram $h$, $|\mathrm{orbit}(h)|\in\{2,4\}$.
@@ 2.3 Hamming Distances of Group Actions
Lemma 2.11 (Key Lemma).
If $d_{H}(h,\mathrm{rev}(h))=6$, then $\mathrm{rev}(h)=\mathrm{comp}(h)$.
@@
If all 6 positions differ between $h$ and $\mathrm{rev}(h)$, then for each $i$: $\mathrm{rev}(h)(i)=h(5-i)\neq h(i)$. Since bits are in $\{0,1\}$, this means $h(5-i)=1-h(i)=\mathrm{comp}(h)(i)$. Thus $\mathrm{rev}(h)=\mathrm{comp}(h)$. ∎
This lemma is crucial: when reversal achieves maximum distance 6, it coincides with complement, so there is no choice to make.
@@ 4.2 Cost Minimization
On a size-4 orbit, a third equivariant pairing exists: $\{h,\mathrm{comp}(\mathrm{rev}(h))\}$. For orbits where $d_{H}(h,\mathrm{rev}(h))=4$, we have $d_{H}(h,\mathrm{comp}(\mathrm{rev}(h)))=2$, making the $\mathrm{comp}\circ\mathrm{rev}$ pairing strictly cheaper. A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits achieves total cost $96<120$. However, this requires case-by-case analysis of each orbit, whereas the reverse-priority rule is a uniform principle. The King Wen sequence follows the uniform rule rather than the per-orbit optimum.
@@ 5.2 Uniqueness Mechanism
The key to uniqueness is Lemma 2.11: when $d_{H}(h,\mathrm{rev}(h))=6$, we have $\mathrm{rev}(h)=\mathrm{comp}(h)$, so the two candidate partners coincide. This eliminates all apparent degrees of freedom in the optimization.

### v2

@@ 1 Introduction
The Boolean hypercube $\{0,1\}^{n}=\{0,1\}^{n}$ admits two natural involutions: bitwise complement $\mathrm{comp}$, which flips each bit ($0\leftrightarrow 1$), and bit reversal $\mathrm{rev}$, which reverses bit order ($b_{0}b_{1}\cdots b_{n-1}\mapsto b_{n-1}\cdots b_{1}b_{0}$). Together these generate the Klein four-group $K_{4}=\{\mathrm{id},\mathrm{comp},\mathrm{rev},\mathrm{comp}\circ\mathrm{rev}\}$.
A $\mathrm{comp}/\mathrm{rev}$ matching is a perfect matching where each pair $\{h,h^{\prime}\}$ satisfies either $h^{\prime}=\mathrm{comp}(h)$ or $h^{\prime}=\mathrm{rev}(h)$. More generally, a $K_{4}$-equivariant matching also permits $h^{\prime}=\mathrm{comp}(\mathrm{rev}(h))$.
This paper addresses the optimization problem:
Among $\mathrm{comp}/\mathrm{rev}$ matchings on $\{0,1\}^{n}$, which minimize total Hamming distance (Definition 2.4)?
For $n=6$, one might expect this to be trivial: just pair each element with the cheaper option. But in most equivariant matching problems, locally optimal choices conflict globally—choosing the cheapest partner for $h$ may force a suboptimal choice elsewhere. Our main result is that for $\{0,1\}^{6}$, no such conflicts arise: a simple greedy rule yields the unique global optimum. The key is Lemma 2.12, which shows that $\mathrm{comp}$ and $\mathrm{rev}$ never genuinely compete—whenever they differ in cost, $\mathrm{rev}$ is strictly cheaper.
As an application, we verify that the King Wen sequence—a traditional ordering of the 64 hexagrams of the I Ching into 32 pairs—is isomorphic to this unique optimal matching. We also note that relaxing to full $K_{4}$-equivariant matchings (allowing $\mathrm{comp}\circ\mathrm{rev}$ pairings) yields lower cost (96 vs 120), but requires per-orbit case analysis rather than a uniform rule.
@@ 1.1 Main Results
The King Wen sequence of the I Ching defines a perfect matching on $\{0,1\}^{6}$ that is isomorphic to the unique optimal $K_{4}$-equivariant matching of Theorem $1.$
@@ 2.1 Hexagrams as Binary Vectors
In the I Ching, a hexagram is a figure composed of six stacked horizontal lines, each either solid (representing yang) or broken (representing yin). Reading solid as 11 and broken as 0, from bottom to top, each hexagram corresponds to a 6-bit binary string. There are $2^{6}=64$ hexagrams, corresponding to the vertices of the 6-dimensional hypercube.
@@
A hexagram is an element $h\in\{0,1\}^{6}$, with positions indexed 0 (bottom) through 55 (top).
Remark $2.$
The correspondence between hexagrams and binary numbers was noted by Leibniz [2].
@@ 2.2 The Klein Four-Group Action
For any hexagram $h$, $|\mathrm{orbit}(h)|\in\{2,4\}$ (see Figure 1).
@@
[…]
Figure 1: Orbit types under $K_{4}$. Edge labels show Hamming distances. In (a), the $\mathrm{rev}$ edge (blue) has cost 2244, always $\leq$ the $\mathrm{comp}$ edge (cost 66). In (b) and (c), only one non-trivial pairing exists.
@@ 2.3 Hamming Distances of Group Actions
Lemma 2.12 (No-Conflict Lemma).
For any hexagram $h$, exactly one of the following holds:
  1. 1.
    $d_{H}(h,\mathrm{rev}(h))<d_{H}(h,\mathrm{comp}(h))=6$ (reversal is strictly cheaper), or
  2. 2.
    $\mathrm{rev}(h)=\mathrm{comp}(h)$ and both distances equal 66 (the options coincide), or
  3. 3.
    $\mathrm{rev}(h)=h$ (palindrome, so $\mathrm{rev}$ yields trivial pairing).
In particular, $\mathrm{rev}$ and $\mathrm{comp}$ never genuinely compete: when both are valid and distinct, $\mathrm{rev}$ is strictly cheaper.
@@
Since $\mathrm{comp}$ flips all bits, $d_{H}(h,\mathrm{comp}(h))=6$ always. For $\mathrm{rev}$: if $d_{H}(h,\mathrm{rev}(h))=6$, then every position differs, so $h(5-i)\neq h(i)$ for all ii. This means $h(5-i)=1-h(i)=\mathrm{comp}(h)(i)$, hence $\mathrm{rev}(h)=\mathrm{comp}(h)$. If $\mathrm{rev}(h)=h$, then $h$ is a palindrome. Otherwise $d_{H}(h,\mathrm{rev}(h))<6$. ∎
Remark 2.13 (Why Greedy Works).
In typical equivariant matching problems, choosing the locally optimal partner for $h$ may force suboptimal choices elsewhere. The No-Conflict Lemma eliminates this tension: the greedy choice (prefer $\mathrm{rev}$ unless it’s trivial) is unambiguous. There is no trade-off to optimize, hence no room for conflict (see Figure 2). This is why the simple reverse-priority rule achieves the unique global optimum.
[…]
Figure 2: Forced consistency of the reverse-priority rule. Choosing $\mathrm{rev}(h)$ as partner for $h$ simultaneously determines $h$ as partner for $\mathrm{rev}(h)$, since $\mathrm{rev}$ is an involution. No secondary decision can conflict.
@@ 4.2 Cost Minimization
On a size-4 orbit, a third equivariant pairing exists: $\{h,\mathrm{comp}(\mathrm{rev}(h))\}$. For orbits where $d_{H}(h,\mathrm{rev}(h))=4$, we have $d_{H}(h,\mathrm{comp}(\mathrm{rev}(h)))=2$, making the $\mathrm{comp}\circ\mathrm{rev}$ pairing strictly cheaper (see Figure 3). A mixed strategy using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits achieves total cost $96<120$. However, this requires case-by-case analysis of each orbit, whereas the reverse-priority rule is a uniform principle. The King Wen sequence follows the uniform rule rather than the per-orbit optimum.
[…]
Figure 3: A size-4 orbit where $\mathrm{comp}\circ\mathrm{rev}$ is optimal. The diagonal pairing (dashed) has cost 22, versus 44 for reversal (solid). Using $\mathrm{comp}\circ\mathrm{rev}$ on such orbits reduces total cost from $120$ to $96$, but requires per-orbit analysis.
@@ 5.2 Uniqueness Mechanism
The key to uniqueness is the No-Conflict Lemma 2.12: $\mathrm{comp}$ and $\mathrm{rev}$ never genuinely compete. When they have different costs, $\mathrm{rev}$ wins; when they have equal cost, they coincide. This eliminates all apparent degrees of freedom, making greedy optimization globally consistent.